Circular No. 5654

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 H. W. Duerbeck and E. K. Grebel, European Southern Observatory,

and D. Beele, Astronomisches Institut Münster, report: “Spectrograms (360-1000 nm) taken Nov. 15.26 UT with the ESO/MPI 2.2-m telescope (+ EFOSC II) at La Silla indicate that the spectral appearance has changed dramatically in two days’ time (cf. IAUC 5653). The nova reached maximum on Nov. 13 and now shows a rich emission line spectrum of H, N I, O I, Na I, Ca II, and Fe II (multiplets 27, 37, 38, 42, 49, 73, 74). It also shows diffuse enhanced features: weak absorption components of H yield expansion velocities of –1100, –2630, and possibly –4050 km/s. A rapid brightness decline is indicated in Strömgren photometry obtained with the Danish 0.5-m telescope at ESO: Nov. 12.20, y = 10.37; 13.24, 10.20; 15.17, 11.00; 16.16, 11.60.”

PERIODIC COMET SWIFT-TUTTLE (1992t) B. Jones and R. K. Pina, University of California at San Diego; and

E. F. Milone, Rothney Astrophysical Observatory, University of Calgary, report that 11.7-μm images (0′′.83/pixel) of P/Swift-Tuttle were obtained on Nov. 7.1 and 8.1 UT by Pina, J. R. Sarmecanic, and Milone (field-of-view of 30′′ × 17′′) and on Nov. 12.1 and 13.1 by R. C. Puetter and J. McLaughlin (19′′ × 24′′), using the UCSD 10-μm array camera at the Mt. Lemmon Infrared Observatory. The comet showed a decrease in flux of 28 percent between the two dates, going from 37.7 to 27.3 Jy in a 5′′.8 beam centered on the nucleus.

R. D. Gehrz, G. Lawrence, and K. Flugaur, University of Minnesota; and C. E. Woodward, University of Wyoming, report thermal infrared measurements of P/Swift-Tuttle on Nov. 12.0 UT using a multifilter GaGe bolometer with a 9′′ diaphragm on the 0.76-m telescope of O’Brien Observatory. The chopper throw placed the reference beams 45′′ north and south of the coma center. The infrared broadband magnitudes are K [6.53 (3σ upper limit), L = 5.9 ± 0.3, M = 3.0 ± 0.3, N (7-14 μm) = –0.2 ± 0.1, Q (18-22 μm) [–2.0 (3σ upper limit). Additional photometry through the narrow-band IRTF silicate filter set gave magnitudes of [7.8 μm] = 0.7 ± 0.3, [8.7 μm] = –0.3 ± 0.1, [9.8 μm] = –0.7 ± 0.15, [10.3 μm] = –0.64 ± 0.3, [11.6 μm] = –1.4 ± 0.3, and [12.5 μm] = –1.1 ± 0.1. The data are consistent with continuum emission from a 367-K blackbody. There is no evidence for a strong 10-μm silicate emission feature.

1992 November 16 Daniel W. E. Green